If the probability density function of a random variable X is given byย 

๐‘“(๐‘ฅ) = {2๐‘˜๐‘ฅ๐‘’

โˆ’

๐‘ฅ2

, ๐‘ฅ > 0

0,

๐‘ฅ

โ‰ค


Determine (i)k (ii)distribution function.


1
Expert's answer
2022-02-23T10:17:44-0500
"f(x)= \begin{cases}
 2kxe^{-x^2} &x>0\\
 0 &x\le0
\end{cases}"

(a)


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=1"

"\displaystyle\int_{0}^{\infin} 2kxe^{-x^2}dx=\lim\limits_{t\to \infin}\displaystyle\int_{0}^{t} 2kxe^{-x^2}dx"

"=k\lim\limits_{t\to \infin}[-e^{-x^2}]\begin{matrix}
 t \\
 0
\end{matrix}=k=1"

"k=1"

"f(x)= \begin{cases}
 2xe^{-x^2} &x>0\\
 0 &x\le0
\end{cases}"

(b)


"F(x)=\displaystyle\int_{-\infin}^{x}f(y)dy"

If "x\le0, F(x)=0."

If "x>0"

"F(x)=\displaystyle\int_{0}^{x}2ye^{-y^2} dy=[-e^{-y^2}]\begin{matrix}
 x \\
 0
\end{matrix}=1-e^{-x^2}"

"F(x)= \begin{cases}
 0 &x\le0\\
 1-e^{-x^2} &x>0
\end{cases}"